OpenStudy (sleepyjess):

\(f(x)=(\dfrac{1}{3})^x\) when x=-1

OpenStudy (sleepyjess):

\(-\dfrac{1}{3}\\\dfrac{1}{3}\\3\\-3\)

OpenStudy (sleepyjess):

i think it is \(-\dfrac{1}{3}\) but im not sure

hartnn (hartnn):